### v1

# Poisson Hail on a Wireless Ground

Abstract
This paper defines a new model which incorporates three key ingredients of a large class of wireless communication systems: (1) spatial interactions through interference, (2) dynamics of the queueing type, with users joining and leaving, and (3) carrier sensing and collision avoidance as used in, e.g., WiFi. In systems using (3), rather than directly accessing the shared resources upon arrival, a customer is considerate and waits to access them until nearby users in service have left. This new model can be seen as a missing piece of a larger puzzle that contains such dynamics as spatial birth-and-death processes, the Poisson-Hail model, and wireless dynamics as key other pieces. It is shown that, under natural assumptions, this model can be represented as a Markov process on the space of counting measures. The main results are then two-fold. The first is on the shape of the stability region and, more precisely, on the characterization of the critical value of the arrival rate that separates stability from instability. The second is of a more qualitative or perhaps even ethical nature. There is evidence that for natural values of the system parameters, the implementation of sensing and collision avoidance stabilizes a system that would be unstable if immediate access to the shared resources would be granted. In other words, for these parameters, renouncing greedy access makes sharing sustainable, whereas indulging in greedy access kills the system.
1 Taxonomy and Motivations
We discuss here four basic classes of spatial queuing dynamics. In each case, there is a time-space Poisson process of customer arrivals on a compact ${\cal X}\subset\mathbb{R}^{d}$. Arrivals are characterized by a time $t\in\mathbb{R}$ and a locus $x\in\cal X$. Each arrival also comes with a service demand (or file size) $h\in{\mathbb{R}}^{+}$, that may vary from customer to customer. The compact $\cal X$ can be seen as the server. The intensity measure of this Poisson process is assumed to be of the form $\lambda\mathrm{d}t\mu(\mathrm{d}x)$, where $\mu$ is a measure on $\mathcal{X}$ with $\mu(\cal X)<\infty$.
It is best to start with the Poisson Hail on a Hot Ground (PHHG) dynamics introduced in [1]. In this model, in addition to their service demand, each arrival at locus $x$ comes with an exclusion set $S\subset\mathbb{R}^{d}$. A typical instance, which is used in this introduction, is that where $S=B(x,R)$, namely the closed ball with radius $R$ centered at $x$, where $R>0$ is a fixed constant or a random variable. Two arrivals are said to be conflicting if their exclusion sets intersect. No two conflicting customers can be served simultaneously. The service of a customer with arrival time $t$ starts when all conflicting customers that arrived before $t$ have left the system. This can be seen as a local First-Come-First-Serve (FCFS) rule. The service speed is constant. Here is what explains the terminology. The ball of a customer is the footprint of the hailstone associated with this customer and its service demand is the height of this hailstone. The temperature of the compact $\cal X$ (subset of the ground) is constant so that the hailstones touching the ground melt at constant speed, whereas the others form a (cold) heap-shaped spatial queue.
The second dynamics is that obtained from the PHHG when setting $S=\{x\}$ (or $R=0$), namely a hailstone arriving at $x$ is an ice-stick at $x$. In this case, when the space marginal measure $\mu$ is absolutely continuous w.r.t. Lebesgue measure, no two customers ever conflict a.s., so that each stick receives immediate attention and melts at constant speed. This model is the spatial birth and death (SBD) process introduced in [2]. If the space marginal is atomic, we get a queuing network made of a finite collection of independent M/M/1 queues located at the atoms of $\mu$.
The third model is the wireless spatial birth and death (WSBD) process which is a variant of the SBD introduced in [3]. The only difference with the SBD is in the service rate (or melting speed) of sticks. Rather than constant speed, the service rate of a customer at $x$ is proportional to the Shannon rate at $x$ when considering all other concurrent customers as interferers and treating interference as noise. The rationale for this service rate comes from information theory and wireless communications [4]. The service discipline is hence immediate attention but with varying service rates across time and space.
The present paper is focused on a model which is new to the best of our knowledge and that we will refer to as the Poisson hail on a wireless ground (PHWG) model. It is meant to study arrivals with (non-trivial) exclusion sets and varying service rates across time and space. In the wireless setting, the exclusion set serves as a protection zone around wireless transmissions, which is a way to control the medium access similar to what is introduced in carrier-sense multiple access protocols (see below). The dynamics can be described as follows:
  • Local FCFS discipline: A hailstone only starts being served (being active) when all conflicting hailstones that precede it have left.
  • Shannon rate: The service rate of each active hailstone at any time $t$ is its Shannon rate, which is a deterministic function of the configuration of all active hailstones at time $t$ (see Eq (1)). Note that at a given time $t$, this rate may be different for two customers being served at that time.
  • Departure at service completion: As soon as a hailstone has fully melted (its service demand is exhausted), it leaves the system.
This model can be seen as a variant of previous models in two equivalent ways:
  • The variant of the PHHG model when the temperature of the ground varies with time and space according to Shannon’s capacity formula, namely the melting speed at time $t$ of a hailstone touching the ground and centered at $x$ is given by the Shannon rate, which depends on all hailstones touching the ground at that time.
  • The variant of the WSBD model where, rather than starting transmitting immediately upon arrival, an arrival is required to apply certain priority rules based on their exclusion sets and which are those of the PHHG model.
The rationale for introducing this model goes beyond the sake of completeness. As suggested by (parts of) the terminology, the practical motivations are to be found in the analysis of wireless communications (see Section 2) at the system level. System-level properties include, for instance, the capacity region of the queuing system of the mean system time of a typical customer. By capacity, we mean the queuing theoretic capacity, namely the set of values for $\lambda$ such that the dynamics in question is stable (say positive recurrent under a proper Markov representation) when fixing all other system parameters. The fact that this stability region of PHWG is of the form $\lambda<\lambda_{c}$ when all other parameters are fixed is one of the main results established in the present paper, together with a constructive representation of $\lambda_{c}$. As we shall see, there are system parameter regions for which the WSBD is unstable, whereas the PHWG is stable. In more concrete terms, in the stability region in question, the queuing delays incurred by some customers due to the exclusion/protection rule in fact allow one to get finite system times in the steady state, whereas stationary system times are infinite in the situation with immediate access to the channel ($R=0$). We find it interesting that refraining from immediate access to a system could stabilize the system in question. This can be seen as a strong argument in favor of protocols of the CSMA-CA-type (see Section 2), which control the access to the shared medium, in order to prevent that nearby customers access that shared channel at the same time. In the stable case, we also analyze the stationary distribution of the waiting or system time of a typical customer and its dependence on $R$. Even if wireless motivations are essential, we will show (see Section 6) that this class of models has, in fact, a much broader scope and potential set of applications.
2 Contributions
The SBD dynamics was first studied in [2] and was then the object of numerous further studies. The PHHG dynamics was introduced in [1]. The stability condition is known. The steady state can be constructed for both compact and infinite spaces. Both dynamics are rather well understood. The WSBD model was introduced in [3] for the compact and continuous space case. The infinite discrete space case was analyzed in [5], where the stability conditions were obtained and a closed form for the steady state queue size was derived. The infinite continuous state space case is unsolved at this stage, to the best of our knowledge.
As explained above, the PHWG model, which is the central topic of the present paper, admits both WSBD and PHHG (and hence SBS) as special cases. In view of the state of the art for these special cases, we will restrict ourselves to the compact state space case. The first objectives are the construction of the steady state and the identification of the stability region of PHWG. Our main result on the matter is the proof of the fact that, when fixing all parameters of the model except $\lambda$, there exists a critical value $\lambda_{c}$ such that if $\lambda<\lambda_{c}$, the PHWG dynamics is a positive recurrent continuous Markov chain on the space of counting measures, whereas it is transient if $\lambda>\lambda_{c}$. Another aim is the comparison of PHWG with all the dynamics listed above. One specific example is the identification of the parameter region where (in some sense to be defined more precisely below) PHWG does better than WSBD. Since WSBD is obtained when setting $R=0$ in PHWG, a simple instance is the identification of the set of parameters other than $R$ such that there is an optimum of $\lambda_{c}$ for $R=R^{*}$ with $R^{*}\neq 0$, and the determination of $R^{*}$. A third aim of the paper is the extension of this model to spatial birth and death with interaction (or death) functions other than that associated with the Shannon rate. As already mentioned, we show that there is a wide class of dynamics where the approach and results developed here extend in a natural way.
2.1 Mathematical Approach
A key step in the construction of the steady state and the identification of the stability region is the identification of arrival blocks that form i.i.d sequences. A sufficient condition to construct such blocks is the existence of what we call zigzag events in the exclusion sets. The simplest such event consists in a single exclusion set covering the whole space. As we show below, more refined zigzag events occur whenever each exclusion set contains a ball of deterministic radius $R$ where $R$ is any positive constant common to all arrivals. More refined events can be considered, linked to the rank of tropical matrices.
2.2 Implications for Wireless Networks
The main novelty of the paper in terms of wireless communications is in the analysis of the dynamics of the spatial properties of carrier-sense multiple access with collision avoidance (CSMA/CA). The principle of CSMA/CA consists in sensing the shared medium and in refraining accessing it if there is a nearby transmitter already using it. There is an immense literature on this multiple access scheme. Its spatial but static analysis was first considered using a Matérn point process heuristic in [6], without analysis of dynamic stability. [7] defines and analyzes the conditions for $\epsilon$-stability, where at most $\epsilon$-fraction of wireless queues are instable. Mean-field heuristics of the dynamics were proposed [8], with an approach ignoring the spatial nature of the problem and in [8] when taking this feature into account. To the best of our knowledge, the present paper is the first to propose an exact (no mean-field approximation) analysis of the spatial properties of CSMA/CA type dynamics.
Let us stress that the version of the collision avoidance mechanism studied in the present paper is quite different from that used in the current versions of the CSMA-CA protocol. The main common feature is the implementation of an exclusion principle which forbids that two (or more customers) with a mutual distance smaller than a given value could transmit simultaneously. The FCSF mechanism used in the paper in order to implement this is quite different from what is currently done.
2.3 Paper Organization
The paper is structured as follows: Section 3 defines the dynamics. Section 4 establishes the main stability result. Section 5 gives simulation and comparison results. Section 6 gives the class of Spatial Interaction Birth and Death (SIBD) processes to which our main results can be extended. Section 7 contains various discussions and a list of open problems. Finally, the appendix gathers the proofs of our two main theorems.
3 Description of the Dynamics
3.1 Arrival Process
We model the arrival of customers as a stationary marked Poisson space-time process of intensity $\lambda$ on $\cal X\times\mathbb{R}$, denoted by $\mathcal{A}=\{\left((x_{i},t_{i}),(S_{i},h_{i})\right)\}_{i\in\mathbb{N}^{+}}$, where $(x_{i},t_{i})$ denotes the location and arrival time of the $i$-th customer, ordered by increasing time of arrival. Each arrival has a mark (or “hailstone”) $(S_{i},h_{i})$, where $S_{i}\subset{\cal X}$ is the “footprint” (e.g. a ball) and $h_{i}\in\mathbb{R}^{+}$ is the “height” of the hailstone. The random sequence of marks $\{(S_{i},h_{i})\}$ is assumed to be i.i.d. and not to depend on $\{(x_{i},t_{i})\}$. $S_{i}$ and $h_{i}$ need not be independent. In what follows, we omit the index $i$ if it is not needed. For instance, $S$ is a random variable with the same distribution as $S_{1}$. We assume that $h$ has a finite mean. In the wireless setting, the $i$-th customer is a wireless link where the transmitter and receiver are co-located. This simplification, which makes sense when the distance from transmitter to receiver is small compared to the distances between links, is that used in [3].
The dynamics is depicted in Figure 1. A discrete-time version is illustrated for simplicity.
Figure 1: A discrete-time illustration of the dynamics on $\cal X\subset\mathbb{R}$. The hot ground (or equivalently, $\cal X$) is illustrated by the red line segment. Each rectangle represents an arrival, where the height represents the file size, and the projection of the rectangle on the ground represents the exclusion set.
3.2 Service Discipline
As already explained, the service discipline is locally FCFS, i.e., the service of customer $n$ (arriving at time $t_{n}$ at $x_{n}$) starts as soon as all customers $k$ arriving before $t_{n}$ and such that $S_{n}\cap S_{k}\neq\emptyset$ have left. The attenuation function $l$, a function from $\mathbb{R}^{+}$ to $\mathbb{R}^{+}$, is assumed to be non-increasing and bounded from above. The positive constants $b,s$, and $w$ represent the bandwidth, the signal power and the additive white Gaussian noise power, respectively. With this notation, the service rate of customer $n$ at any time $t$, when it is active and still in the system, is
$b\log\left(1+\frac{s}{w+\sum_{x_{k}\in\phi_{t}\setminus\{x_{n}\}}l(||x_{k}-x_{n}||)}\right),$ (1)
with $\phi_{t}$ the point process made of all points $x_{k}$ of customers still in the system and being currently served at time $t$ (this excludes those customers which have precedence constraints w.r.t. other customers present in the system at that time).
3.3 Construction of model in $\mathbb{R}^{2}$ with Independent File Size
For this first description, we focus on the case with state space in $\mathbb{R}^{2}$, with Poisson space-time arrivals with intensity $\lambda$ and with service demands (file sizes) which are i.i.d. with distribution $F$ with positive and finite mean. The compact phase space is assumed to be $[-W,W]^{2}$. Its torus version is used. Concerning exclusion sets, a special case of interest is that where the exclusion set of a customer arriving at $x$ contains a ball centered at $x$ and with fixed radius $r>0$.
We define a sequence of directed acyclic graphs (DAGs)  [^4] A DAG is a directed graph with no directed cycles., denoted by $\{G_{n}\}$, $n=0,1,2,\ldots,$ and an increasing sequence of times $\{j_{n}\}$ by induction w.r.t. $n$. The continuous-time process that we build is piecewise constant with jump times $\{j_{n}\}$, where $j_{n}$ is a strictly increasing sequence. DAG $G_{n}$ is the state DAG on the time interval $[j_{n},j_{n+1})$. A DAG is possibly non-connected. DAG $G_{n}$ features a set of vertices (customers) $V_{n}$ with acyclic directed edges $E_{n}$ between them. The service (or active) set $U_{n}$ of $G_{n}$ is the subset of vertices of $V_{n}$ without predecessor [^5] Vertex $v$ is a predecessor of $u$ if there exists a directed edge from $v$ to $u$. . Each node $v\in V_{n}$ has a location $x(v)\in[-W,W]^{2}$, an exclusion set $S(v)$, a birth time $b(v)$, and a death time $d(v)$. The vertices of $G_{n}$ have each an additional state variable which is its residual service demand (file size) at time $j_{n}$, denoted by $z_{n}(v)\in{\mathbb{R}}^{+}$.
In order to define the dynamics, we need
  • an i.i.d. sequence of exponential random variable $\{\widetilde{\tau}_{n}\}$ of parameter $\lambda 4W^{2}$; $\widetilde{\tau}$ represents the time that elapses between $j_{n}$ and the next potential arrival after time $j_{n}$.
  • an i.i.d. sequence of uniform random variables $\{\widetilde{X}_{n}\}$ on $[-W,W]^{2}$; $\widetilde{X}_{n}$ represents the locus of the potential next arrival after time $j_{n}$.
  • an i.i.d. sequence of random variables $\{\widetilde{S}_{n}\}$; $\widetilde{S}_{n}$ represents the footprint of the next potential arrival after time $j_{n}$.
  • an i.i.d. sequence of random variables $\{\widetilde{h}_{n}\}$ of distribution $F$; $\widetilde{h}_{n}$ represents the file size of the potential next arrival after time $j_{n}$.
The word “potential” here means that this variable can be used or discarded depending on conditions spelled out in the induction. The tilda is used to distinguish it from the effective variable. The induction is as follows. When the DAG at $j_{n}$ is $G_{n}$, define the potential departure time of node $v\in U_{n}$ as
$p_{n}(v)=z_{n}(v)/r_{n}(v),$
with $r_{n}(v)$ the Shannon rate of $v$ in the spatial configuration
$\Psi_{n}:=\{{v\in U_{n}}\colon x(v)\},$ (2)
namely
$r_{n}(v)=b\log\left(1+\frac{s}{I_{n}(v)+w}\right),$ (3)
with
$I_{n}(v)=\sum_{u\in\Psi_{n}\setminus\{v\}}l(||v-u||).$ (4)
Let also
$q_{n}:=\min_{v\in U_{n}}p_{n}(v).$ (5)
  • If $q_{n}>\widetilde{\tau}_{n}$, then
    • $j_{n+1}:=j_{n}+\widetilde{\tau}_{n}$.
    • $V_{n+1}:=V_{n}\cup\{v^{\prime}\}$ where the extra node $v^{\prime}$ is a new arrival which is allocated location $\widetilde{X}_{n+1}$, exclusion set $\widetilde{S}_{n+1}$, and residual service time $\widetilde{h}_{n+1}$, i.e., $x(v^{\prime})=\widetilde{X}_{n+1},~{}b(v^{\prime})=j_{n+1},~{}S(v^{\prime})=\widetilde{S}_{n+1},~{}z(v^{\prime})=\widetilde{h}_{n+1}$.
    • $E_{n+1}$ is the union of $E_{n}$ and of the set of edges from ${v^{\prime}}$ to $V_{n}$, defined as follows: there is an edge from $v\in V_{n}$ to ${v^{\prime}}$ if $S(v^{\prime})\cap S(v)\neq\emptyset$. This defines $G_{n+1}$.
    • For all $v\in U_{n}$, $x_{v}$ is unchanged and the residual file size is updated to $z_{n+1}(v)=z_{n}(v)-\widetilde{\tau}_{n}r_{n}(v)$.
    • If ${v^{\prime}}$ has no predecessor in $G_{n+1}$, then $U_{n+1}:=U_{n}\cup\{{v^{\prime}}\}$. Otherwise, $U_{n+1}:=U_{n}$.
  • If $q_{n}{\leq}\widetilde{\tau}_{n}$, then, with $u$ denoting the argmin in the definition of $q_{n}$,
    • $j_{n+1}:=j_{n}+q_{n}$.
    • $V_{n+1}:=V_{n}\setminus\{u\}$ (departure of customer $u$), $d(u)=j_{n+1}$.
    • $E_{n+1}$ is the restriction of $E_{n}$ to $V_{n+1}$.
    • For all $v\in U_{n+1}$, $x(v)$ is unchanged and the residual service time is updated to $z_{n+1}(v)=z_{n}(v)-q_{n}r_{n}(v)$.
    • $U_{n+1}$ is the set of vertices without predecessor in $(V_{n+1},E_{n+1})$. This defines $G_{n+1}$.
This induction step defines $j_{n+1}$, $G_{n+1}$ and the residual service times at time $j_{n+1}$. The initial DAG $G_{0}$ is, for instance, the empty DAG and $j_{0}=0$. Up to ties, which happen with 0 probability, the above dynamics defines a piecewise constant DAG state process up to time $j=\lim_{n\to\infty}j_{n}$. It should be clear that $j=\infty$ a.s. This follows from the fact that $q_{n}$, defined above, is bounded from below by a random variable $Q_{n}$ such that the sequence $\{Q_{n}\}$ is i.i.d. and such that $\mathbb{E}Q_{n}>0$. The latter fact follows from the property that $r_{n}\leq c$ where $c$ is the constant
$c:=b\log(1+\frac{s}{w}).$
Note that each DAG can be represented as a marked counting measure where the atoms of the counting measure are the locations of the customers present in the system and the marks are the precedence edges and the residual service times of these customers (all those having a predecessor in the DAG have their residual file size equal to their full initial file size).
Remark 1.
In the special case where all exclusion sets contain a ball of fixed radius $R>0$, if no two customers with intersecting exclusion sets are served at the same time, the point process of active customers is a hardcore point process with balls of radius $R>0$. This implies that the total number of customers being served at any given time is bounded above by a constant that depends on $W$ and $R$. Since the function $l$ is bounded above, this shows that the interference is bounded above by a constant at any given time and locus. Hence the Shannon rate for any active customer is bounded below by a positive constant at any given time [9].
In the general case, one can still give a positive constant lower bound to the sum of the departure rates of all customers present in the system at any time, given that there is at least one customer of course:
Lemma 2.
Under the foregoing assumptions, if there is at least one customer in the system, then the sum rate of all customers in the system is bounded below by a positive constant.
Proof.
Assuming that there are $N>0$ customers in the system, the sum rate is
$\sum_{n=1}^{N}b\log\left(1+\frac{s}{w+\sum_{x_{k}\in\phi_{t}\setminus\{x_{n}\}}l(||x_{k}-x_{n}||)}\right)$
$\geq$ $b\log\left(1+\sum_{n=1}^{N}\frac{s}{w+\sum_{x_{k}\in\phi_{t}\setminus\{x_{n}\}}l(||x_{k}-x_{n}||)}\right)$
$\geq$ $b\log\left(1+\frac{Ns}{w+(N-1)L}\right),$
where $L$ is the upper bound on $l$. The fact that the function $N\to\frac{Ns}{w+(N-1)L}$ is uniformly bounded below concludes the proof. ∎
3.4 Case with Renewal Arrivals
The previous construction leverages the memoryless property of the exponential distribution. It can easily be extended to the case where the arrival process is a renewal process without multiplicity (rather than a Poisson point process) where the inter-arrival times $t_{i+1}-t_{i}$ are assumed to be i.i.d. and strictly positive with a finite mean, by adding in the induction a supplementary state variable for the remaining time to the next arrival. All steps and results of the last subsection then hold for this case.
4 Stability
This section adopts the assumptions of Section 3. It makes further assumptions on the structure of the exclusion sets.
4.1 Main assumptions on exclusion sets
We present here our set of assumptions.
Let $k\geq 1$ and let $\{B_{i}\}_{i=1}^{k}$ be a finite collection of subsets of ${\cal X}$. Let $B_{0}=\emptyset$. There exists a positive integer $k$ and a sequence of non-empty deterministic measurable sets $B_{1},B_{2},\ldots,B_{k}$ in the space ${\cal X}$ such that
  • (a) $B_{i+1}\cap B_{i}\neq\emptyset$, for any $i=1,2,\ldots,k-1$;
  • (b) $B_{i+1}\setminus\cup_{j=1}^{i}B_{j}\neq\emptyset$, for any $i=1,2,\ldots,k-1$;
  • (c) $\cup_{i=1}^{k}B_{i}={\cal X}$;
  • (d) ${\mathbb{P}}(B_{i}\subseteq S)>0$, for all $i=1,2,\ldots,k$.
Remark 3.
In the special case of $S=\{x\}$ a.s. for an arrival at $x$, the assumptions cannot be jointly satisfied for $|\cal X|$$\neq 0$.
Remark 4.
In the special case where every exclusion set contains a ball of fixed radius $R>0$ centered on the arrival point, it is easy to see that the last property holds for an integer $k$ large enough, which is linked to the number of balls of radius $R$ needed to completely cover the window of interest $\cal X$. However, there may exist a collection of subsets that satisfy the same property for a smaller $k$. For example, the above assumptions hold for $k=1$ when there is a positive probability that an arrival comes with $S=\cal X$. This is the case for i.i.d. exclusion radius with exponential distribution.
Figure 2: An illustration of an underlying deterministic sequence $\{B_{i}\}_{i=1}^{k}$ with the corresponding zigzag and block structure for $\cal X\subset\mathbb{R}$. For simplicity, the arrivals are not illustrated, and here the height of $\{B_{i}\}_{i=1}^{k}$ has no physical meaning.
4.2 Block Representation
Consider the events $A_{i,n}=\{B_{i}\subseteq S_{n}\}$, $n\geq 1$, and, for all $n\geq k+1$, the events
$C_{n}=\left(\bigcap_{i=0}^{k-1}A_{i+1,n-k+i}\right)\bigcap\left(\bigcap_{j=0}^{k-1}A_{k-j,n+j}\right)\equiv C_{n,1}\cap C_{n,2}.$ (6)
In what follows, we will call $C_{n}$ a zigzag event, with $C_{n,1}$ as the “zig” event and $C_{n,2}$ the “zag’ event. We say that $2k$ is the “length” of the zigzag event (i.e. the number of the $A$ events involved), with the lengths of the zig and the zag events to be equal to $k$.
We introduce now some splitting of customers into ”blocks”, or ”cycles”.
Let $n_{0}=0$, $n_{1}=2k\cdot\min\{r\geq 1:{\mathbf{I}}(C_{2kr})=1\}$ and, for $m=1,2,\ldots$,
$n_{m+1}=n_{m}+2k\cdot\min\{r\geq 1:{\mathbf{I}}(C_{n_{m}+2kr})=1\}.$ (7)
Let $T_{0}=0$ and $T_{m}:=t_{n_{m}}$ be the corresponding thinning of the arrival process. For $m\geq 0$, let
$Y_{m}:=\{((x_{n},t_{n}-t_{n_{m}}),(S_{n},h_{n})),n=n_{m}+1,n_{m}+1,n_{m}+2,\ldots,n_{m+1}\}.$
Definition 5.
For $m\geq 1$, the customers numbered $n_{m},n_{m}+1,\ldots,n_{m+1}-1$ form the $m$-th block of customers.
  • The first customer of this block (that having an exclusion set containing $B_{k}$) is the opening customer of block $m$; it arrives at $T_{m}$.
  • The last customer of this block (which also has an exclusion set containing $B_{k}$) is the closing customer of block $m$; it arrives at $\hat{T}_{m}:=t_{n_{m+1}-1}$.
Note that these blocks have independent but not identically distributed components. For example, the $k$ first arrivals of a block have exclusion sets $S_{1},\ldots,S_{k}$ such that $B_{k-i+1}\subseteq S_{i}$ for all $i=1,\ldots,k$.
Fig. 2 illustrates a zigzag and block structure based on a deterministic set of $\{B_{i}\}_{i=1}^{k}$ on $\cal X\subset\mathbb{R}$. The sequence of arrivals has exclusion sets that cover the sequence $\{B_{i}\}_{i=1}^{k}$ and are omitted. Such sets can easily be constructed for systems in $\mathbb{R}^{d}.$ One can see that the blocks are formed by “zag-zig” events and not by “zig-zag” events.
The proof of the next proposition follows directly from the strong Markov property.
Proposition 6.
The sequence of random elements $\{Y_{m}\}_{m\geq 1}$ is i.i.d. and does not depend on $Y_{0}$.
Remark 7.
As already mentioned, there are several other options for defining zigzag events while preserving the properties identified above and below. One of them is that with no-doubling the last event in the zig. For example, if $k=1$ (i.e., we have customers with exclusion set equal to the whole space), then we do not have to double this event and all the properties identified still hold. The specific structure proposed here is retained for its simplicity and symmetry. This concept comes from tropical mathematics and is linked to what is now called the tropical matrix rank. See, e.g., [10].
Note also that, in Subsection 4.1, we may add for free a further assumption (e) to assumptions (a)–(d):
  • (e) ${\mathbb{P}}(B_{i}\cup B_{j}\subseteq S)=0$, for all $i\neq j$.
Indeed if (e) fails for some  $i\neq j$, then we may merge these two sets and consider a new collection of $(k-1)$ events $B$’s instead.
There are also other options for defining the blocks. For example, we may deal with smaller blocks, by defining $\{n_{m}\}$ inductively as
$n_{m+1}=\min\{n\geq n_{m}+2k:\ {\mathbf{I}}(C_{n})=1\}.$ (8)
However, we prefer the definition (7) because the variant with (8) leads to a more complicated construction.
The index of the next block starts as early as $n_{m}+2k$.
4.3 Reduction to a Lindley type recurrence relation
In view of the service discipline and the definition of blocks, customers are served block by block, i.e., the service of the customers of block $m+1$ may start only when the service of the closing customer of block $m$ is completed, and at that time, all customers of block $m$ have already left. This can be exploited to derive a Lindley-type evolution equation.
For any customer, define its waiting time as the time that elapses between its arrival time and the epoch when it starts its service. Let $D_{m}$ denote the departure time of the closing customer of block $m$. Thanks to the block-by-block service property, the waiting time $W_{m+1}$ of the opening customer of block $m+1$ is equal to $(D_{m}-T_{m+1})^{+}$.
We have $D_{m}=T_{m}+W_{m}+\sigma_{m}$, where $\sigma_{m}$ is the time to serve all customers of block $m$ (or equivalently the time to complete the service time of the closing customer of this block) when its opening customer starts its service at time $T_{m}+W_{m}$. The key observation now follows from the construction by induction of Section 3, is that the random variable $\sigma_{m}$ is a measurable function of $W_{m}$ and $Y_{m}$:
$\sigma_{m}=\Sigma(W_{m},Y_{m}).$
Proposition 8.
The random variables $W_{n}$ satisfy the recurrence relation
$W_{m+1}=\left(W_{m}+\Sigma(W_{m},Y_{m})-(T_{m+1}-T_{m})\right)^{+}.$ (9)
Proof.
We have
$W_{m+1}$ $=(D_{m}-T_{m+1})^{+}$
$=(T_{m}+W_{m}+\sigma_{m}-T_{m+1})^{+}$
$=\left(W_{m}+\sigma_{m}-(T_{m+1}-T_{m})\right)^{+}$
$=\left(W_{m}+\Sigma(W_{m},Y_{m})-(T_{m+1}-T_{m})\right)^{+}.$
Note that $W_{m}$ is a function of $\{Y_{p}\}_{p<m}$ whereas conditionally on $W_{m}=W$, $\Sigma(W_{m},Y_{m})$ and $T_{m+1}-T_{m}$ are measurable functions of $Y_{m}$. Hence the dynamics of the sequence $\{W_{m}\}$ admits the representation
$W_{m+1}=f(W_{m},Y_{m}).$
Therefore,
Proposition 9.
The sequence $\{W_{m}\}$ forms a time-homogeneous Markov chain.
Let
$\widehat{\sigma}_{n}=\Sigma(\infty,Y_{n}).$
Note that the last random variable can be interpreted as follows: Think about all arrivals of block $Y_{n}$ to be already available at time 0. These customers then form a heap of hailstones. Start serving all customers of this heap following the induction steps of Section 3, namely in the order prescribed by the precedence relations between these customers and at any time, at the individual rate prescribed by the induction steps. Then $\widehat{\sigma}_{n}$ is the epoch at which the last element of this heap leaves.
Furthermore, since the random variables $\{Y_{n}\}$ are independent for $n\geq 1$ and identically distributed for $n\geq 2$, the same holds for $\widehat{\sigma}_{n}$: they are independent for $n\geq 1$ and identically distributed for $n\geq 2$.
Lemma 10.
Under the foregoing assumptions, $\mathbb{E}\widehat{\sigma}_{m}<\infty,$ for any $m$. Moreover, there exists an absolute bound, say $K$, such that ${\mathbb{E}}\sigma_{m}(x)\leq K$, for all $x$ and $m$. In addition, the random variables $\{\sigma_{m}(x),m\geq 1,x\geq 0\}$ are uniformly integrable and, for any $m$, we have $\sigma_{m}(x)\to\widehat{\sigma}_{m}$ a.s., as $x\to\infty$. Therefore, for all $m$, ${\mathbb{E}}\sigma_{m}(x)\to{\mathbb{E}}\widehat{\sigma}_{1}$ as well.
Proof.
The number of customers in any block, say
$\nu_{m}:=n_{m}-n_{m-1}$
(see (7)), has a light-tailed distribution since it is nothing else than a geometrically distributed random variable with parameter
$p_{b}:={\mathbb{P}}(C_{n})=\left(\prod_{i=1}^{k}{\mathbb{P}}(B_{i}\subseteq S)\right)^{2}>0$
multiplied by $2k$. Further, since the total service rate (of all customers) in a non-empty system is bounded below by a positive constant, say $\kappa$ (see Remark 1), the total time for service of all customers in a block cannot exceed the sum of their heights, divided by $\kappa$. Note also that the random variables
$\sum_{i=n_{m}-k}^{n_{m}-1}h_{i},\ \ m=1,2,\ldots$
are i.i.d. Then, thanks to the Wald identity, we have that the mean total height of all customers in the block number $m=2,3,\ldots$ is equal to
${\mathbb{E}}\sum_{j=0}^{\nu_{m}-1}\sum_{r=0}^{2k-1}h_{n_{m-1}+2kj+r}$ $={\mathbb{E}}\sum_{j=1}^{\nu_{m}}\sum_{r=-k}^{k-1}h_{n_{m-1}+2kj+r}$
$={\mathbb{E}}\nu_{m}{\mathbb{E}}\sum_{r=0}^{2k-1}h_{r}=2k{\mathbb{E}}\nu_{1}{\mathbb{E}}h_{1}=:R_{1}<\infty.$
In the first block, only the distribution of the first $k$ random variables $\{h_{i}\}_{i=0}^{k-1}$ differs from that in the other blocks, in the sense that these random variables have the initial (unconditional) distribution. Then, clearly,
${\mathbb{E}}\sum_{j=0}^{\nu_{1}-1}\sum_{r=0}^{2k-1}h_{n_{m-1}+2kj+r}\leq R_{1}+k{\mathbb{E}}h_{1}=:R<\infty.$
Further, for any $m$ and $x$, the value of $\sigma_{m}(x)$ cannot exceed $D_{m}-D_{m-1}$ (recall that $D_{m}$ is the departure time of the last customer in block $m$) and that
$D_{m}-D_{m-1}\leq T_{m}-T_{m-1}+\frac{1}{\kappa}\sum_{j}h_{j}=:Z_{m},$ (10)
where the sum is taken over all customers $j$ from block $m$. Therefore,
${\mathbb{E}}\sigma_{m}(x)\leq{\mathbb{E}}(T_{m}-T_{m-1})+R/\kappa=2k{\mathbb{E}}t_{1}+R/\kappa=:K<\infty.$
Next, the uniform integrability of all $\{\sigma_{m}(x)\}$ follows directly because they all are stochastically bounded by the sum of three random variables that are the same for all $\sigma_{m}(x)$ and that have finite expectations.
Finally, we have the a.s. convergence of $\sigma_{n}(x)$ to $\widehat{\sigma}_{n}$ because, in fact, a stronger “coupling-convergence” takes place:
$\sigma_{m}(x)=\widehat{\sigma}_{m},\ \ \text{ for all}\ \ x>T_{m}-T_{m-1}.$
Indeed, $T_{m-1}+x$ is the time when the last customer in block $m-1$ leaves the system, so if $x>T_{m}-T_{m-1}$, then all customers from block $m$ are already in the system at time $T_{m-1}+x$, like in the saturated system. ∎
4.4 Main Stability Results
We obtain two stability results, one for the Markov chain $\{W_{n}\}$ and one for the continuous-time process $\{W(t),t\geq 0\}$ admitting $\{W_{n}\}$ as an embedded Markov chain.
Below, we use the following notation: $T=T_{1}$, $\nu=\nu_{1}$ (where the latter is defined in the proof of Lemma 10), and $\widehat{\sigma}=\widehat{\sigma}_{1}$. Since $T=\sum_{i=1}^{\nu}t_{i}$, ${\mathbb{E}}\nu=2k/p_{b}$, $\lambda=1/(|{\cal X}|\mathbb{E}t_{1})$ and $\nu$ does not depend on $\{t_{i}\}$, we get
${\mathbb{E}}T=\frac{2k}{\lambda|{\cal X}|p_{b}}.$ (11)
Our first stability result is as follows:
Theorem 1.
Under the foregoing assumptions,
  1. (i)
    If ${\mathbb{E}}T>{\mathbb{E}}\widehat{\sigma}$ – or, equivalently,
    $\lambda<\lambda_{c}:=\frac{2k}{p_{b}|{\cal X}|{\mathbb{E}}\widehat{\sigma}},$ (12)
    then there exists a finite $K>0$ such that the set $[0,K]$ is positive recurrent, i.e., $\tau_{x}\equiv\tau_{x}(K)=\min\{n\geq 1:W_{n}\leq K\}<\infty$ a.s., for any $W_{0}=x$ and, moreover,
    $\sup_{0\leq x\leq K}{\mathbb{E}}\tau_{x}<\infty.$
    In addition, if the distribution of $t_{1}$ has infinite support, then the Markov chain $\{W_{n}\}$ admits a single stationary distribution, say $\pi$, and, for any initial state $x\in{\cal X}$, the distribution of the Markov chain converges to $\pi$ in the total variation norm.
  2. (ii)
    If ${\mathbb{E}}T<{\mathbb{E}}\widehat{\sigma}$ and if there is a constant $\gamma>0$ such that, for any $x$
    ${\mathbb{P}}_{\psi}^{0}(\Sigma(x,Y_{1})>T_{1}+\gamma)\geq\gamma,$ (13)
    then, for some $\varepsilon>0$ and for any initial $x\geq 0$, we have
[…]
    and, in particular, $W_{n}\to\infty$ a.s. as $n\to\infty$.
Remark 11.
In some special cases, there are closed-form expressions for $p_{b}$ and ${\mathbb{E}}\widehat{\sigma}$. For example, for $S_{i}=b(x_{i},R_{i})$ where $\{R_{i}\}$ are exponentially distributed i.i.d. with mean radius $R$, it is sufficient to define the zigzag event based on a single arrival for which the exclusion set covers $\cal X$, that is, $C_{n}=\{{\cal X}\subset S_{n}\}$. In this case, $p_{b}$ is simply the tail probability of an exponential random variable, and we have:
$\lambda_{c}=\frac{\exp(\sqrt{2}W/R)}{|\cal X|{\mathbb{E}}\widehat{\sigma}}.$
The tractability of ${\mathbb{E}}\widehat{\sigma}$ depends both on the distribution of $S_{i}$ and the pathloss function $\ell$. In general, we can estimate $p_{b}$ and ${\mathbb{E}}\widehat{\sigma}$ by Monte Carlo simulation. The details are given in Section 5.
For any time $t\geq 0$, consider an auxiliary system in which we stop arrivals from time $t$ onward and admit only customers arriving before time $t-0$. In this system, there is a (random) finite number of arrivals. Denote by $Z(t)$ the time when the last customer leaves such a system and by $W(t)=Z(t)-t$ the time to empty the system after the interruption of the input process. One can see that $Z(t)$ is finite a.s. (in particular, because the sum of the service rates of customers on the ground is bounded from below) and that $W_{n}=W(T_{n})$, for $n=1,2,\ldots$.
Our second stability result is:
Theorem 2.
The following hold:
  • (i)
    Assume the assumptions of item (i) of the previous theorem hold. If in addition the distribution of $t_{1}$ has a non-lattice distribution, then the Markov process $W(t)$ admits a single stationary distribution and, for any initial finite configuration of customers, the distribution of $W(t)$ converges weakly to its stationary version. If in addition the distribution of $t_{1}$ is spread-out, the convergence holds in the total variation norm.
  • (ii)
    Under the assumptions of item (ii) of the previous theorem, there exists $\delta>0$ such that, for any initial value of $W(0)$,
[…]
The proofs are given in Appendix 8.
Remark 12.
There are three stochastic processes of interest. Two are in discrete-time and one is in continuous-time. The continuous-time process is studied in Theorem 2. It roughly describes the evolution of workload in continuous time. The first discrete-time sequence is the Markov chain studied in Theorem 1. It is obtained when sampling this continuous-time process at discrete epochs which are arrival epochs with an arrival event index $n$ that is a multiple of $2k$ and such that time $t_{n}$ is the center of a zigzag event. There is of course another natural discrete-time stochastic process, which is that of the sequence of workloads at all arrival epochs. It should be clear that, under the stability condition, this last discrete-time sequence admits a unique stationary regime which is reached from all initial conditions. This last stationary regime is obtained from that of the Markov chain of Theorem 1 using the exchange formula of Palm calculus.
5 Simulations
5.1 Estimates of the stability threshold
In view of the expression obtained in (12), the most natural way to estimate $\lambda_{c}$ is to estimate $p_{b}$ and $\mathbb{E}\widehat{\sigma}_{1}$ by Monte Carlo simulation. Both problems can be solved by producing perfect block samples and using them as input to the induction steps of Section 3. For this purpose, in these blocks, there is no need to keep the arrival epochs. The simplest way is to first sample an i.i.d. sequence of arrival loci, height, and exclusion sets until one sees a first $C$ event and to proceed sampling such data until one sees a second $C$ event. This provides of sample of a block by restriction to the arrivals that start with the zag of the first $C$ event and lasts until the zig of the second one. This provides a sample of the number of arrivals in a block, say $N$. One then takes this block as input to the induction steps of Section 3. This in turn provides a perfect sample of $\hat{\sigma}_{1}$. By repeating this, one gets an estimate (together with confidence intervals) of $\mathbb{E}N=1/{p_{b}}$ and of $\mathbb{E}\widehat{\sigma}$.
5.2 Examples
This subsection defines two sets of basic examples, one in the discrete space setting and the other in the continuous space setting.
5.2.1 Discrete-Space Setting
We define here an example in the 1D discrete-space setting. Let $N$ be an even natural number $N\geq 4$. There are $N$ servers located equidistantly on the circle of length $N$. We number them $1$ to $N$ clockwise, say. Customers are represented as rectangles $M\times h$ centered at $x$. Here, $x$ represents the “center” of the customer, which is one of the locations of the servers. Here, $M$ is a fixed odd integer, with $1\leq M\leq N$. The meaning of $M$ is that the customer needs simultaneous service of $M$ neighboring servers. More precisely, the exclusion set of a customer centered at $x$ concerns the $(M-1)/2$ servers on the left and the $(M-1)/2$ servers on the right of server $x$. Exclusion means that if two customers have common servers, then they cannot be served simultaneously. As in the previous discussions, we assume here that customers are served in a locally FCFS way. Finally, $h$ is a positive random variable which represents the file size of a customer.
The distance between two customer centers is the minimum of the two distances between them, clockwise and counterclockwise. Let $d$ be a fixed positive integer. The attenuation function at distance $u\in\mathbb{N}$ is $l(u)=v\cdot 1_{u\leq d}$, where $v>0$ is a constant. The thermal noise is equal to 1. Hence, the service rate of a customer at $x$ and in service is equal to $\frac{1}{1+v\phi_{d}(x)}$ with $\phi_{d}(x)$ with $\phi_{d}(x)$ the number of customers being served at a distance less than or equal to $d$, without counting the customer at $x$.
Note that for $(M-1)/2\geq d$, this model coincides with Poisson Hail since the rate of any customer being served is always 1. It is known that Poisson Hail is strictly monotonic in the footprint of the hailstones. More precisely, for $N$ and $d$ fixed, the value of $\lambda_{c}$ is a decreasing function of $M$, in the range $M\geq 2d+1$.
A special case of this model is obtained by taking $d=N/2$. That if a customer’s service “interferes” with at least another customer, then it interferes with exactly one customer, and its service rate is then $\frac{1}{v+1}$.
If $M>d$, then there is no way to serve two or more customers simultaneously, and we have a single-server queue with service rate 1. If $M=d$, then the simultaneous service of two customers may only occur if the distance between the customer centers is exactly $d$, and at most two customers are served simultaneously.
Consider the saturated system (it means that there are infinitely many customers in the queue in front of the system, which are numbered $1,2,\ldots$) in the case $M=d$. We assume that customer $i$ is centered at $x_{i}$ and brings $H_{i}$ units of work to serve, that $\{x_{i}\}$ is an i.i.d. sequence with elements uniformly distributed on the discrete set of $N$ locations, and that $\{h_{i}\}$ is an i.i.d. sequence of positive random variables.
5.2.2 Continuous-Space Setting
Consider the PHWG model defined on the continuum in 1D. The state space is an interval ${\cal X}=[-W,W]$ of fixed length $2W$. Arrivals follow a space-time Poisson process on ${\cal X}$ with intensity $\lambda$. The ends are glued together to form a circle. In 2D, we consider the state space as a square window ${\cal X}=[-W,W]^{2}$. Arrivals follow a space-time Poisson process on $\cal X$ with intensity $\lambda$. Similarly, the torus version of the model is used to evaluate the wireless scenario.
We consider exclusion sets $\{S_{i}\}$ of two types: (1) balls of random radius centered at the point of arrival, where the random radii are i.i.d. exponential random variables of mean $R$; (2) balls of fixed radius $R$ centered at the point of arrival. The heights $\{h_{i}\}$ are i.i.d. and exponentially distributed.
We simulate the example defined on the torus version of ${\cal X}=[-W,W]^{2}$, $b=1,s=l(0)$. We compare the stability region (captured by the critical threshold) with different fixed or i.i.d. exponentially distributed radius for a fixed function $l$.
We first give the analytical results of two special cases of exclusion sets. For $R\equiv 0$, the model is a WSBD process defined in [3]. The stability region is $\lambda>\lambda_{c}$, where the critical threshold is [3]
$\lambda_{c}=\frac{l(0)}{\ln(2)\mathbb{E}h\int_{\cal X}l(\|x\|)\mathrm{d}x}.$ (14)
On the other hand, the dynamic reduces to that of a M/M/1 queue with global FCFS when the exclusion sets are such that there is at most one customer on the ground at any time. For such an M/M/1 system, the stability region is also known as $\lambda>\lambda_{c}$, and
$\lambda_{c}=\frac{c}{|{\cal X}|\mathbb{E}h},$ (15)
where $c$ is the constant denoting the maximum Shannon rate achieved when there is no interference as defined earlier.
Figure 3: $\cal X$ = $[-2,2]^{2}$. $\mathbb{E}h=1$, i.i.d. exponential radius, and function $l(r)=\min(1,r^{-4}),w=0.05$.
Figure 4: $\cal X$ = $[-10,10]^{2}$. $\mathbb{E}h=1$, fixed constant radius, and function $l(r)=\min(1,r^{-4}),w=0.05$.
In Fig. 3, we consider ${\cal X}=[-2,2]^{2}$, $\mathbb{E}h=1,w=0.05$, and function $l(r)=\min(1,r^{-4})$, with i.i.d. exponential radii. We plot the critical threshold versus the mean radius. From Eq. (14) and (15), the critical thresholds for the special cases $\mathbb{E}R=0$ and $\mathbb{E}R=\infty$ are 0.2558 and 0.2745 for $R=0$ and M/M/1, respectively. The estimation of the stability threshold is done as described in Subsection 5.1 with 1000,000 realizations of blocks (with the exception of the data point at $\mathbb{E}R=0.4$, which is calculated from 3,500 block realizations due to its system complexity).
In Fig. 4, we consider ${\cal X}=[-10,10]^{2}$, $\mathbb{E}h=1,w=0.05$, function $l(r)=\min(1,r^{-4})$, and fixed constant radius. Here, due to the complexity of the system, the stability threshold for the fixed radius at $R\equiv 0.5$ is estimated by comparing the delay distributions of the system with 100,000 arrivals under different arrival intensities.
In both cases, the critical threshold versus exclusion radius displays unimodality: $\lambda_{c}$ has a single peak and the largest value appears in the middle $0<R^{*}<\infty$. This leads to what we refer to as optimal exclusion. The results show that, for the given scenario, moderate protection optimizes the system stability as opposed to immediate access, which is too greedy, or global FIFO, which is too conservative.
6 Spatial Interaction Queuing Processes
This section defines a new and general class of spatial queuing processes. Arrivals in these systems are in terms of e.g. a Poisson rain or a renewal rain as above. Arrivals are independently marked with a locus of arrival $x$ in a given compact $D$ of the Euclidean space, a height $h$ representing the size of the task, and an exclusion set $B$. Service is again locally FCFS, based on the exclusion sets, with the same definition as above. The generalization stems from the fact that the interaction function, which gives the service speed of customers in a given spatial configuration, is general. More precisely, let $\psi$ denote the point process of customers currently being served. For all $x\in\psi$, the service rate of $x$ in this configuration is of the form $s(x,\psi)$, where $s$ is a measurable function from $D\times{\cal M}_{D}$ to $\mathbb{R}^{+}$. Here, ${\cal M}_{D}$ denotes the space of counting measures on $D$. The Shannon rate function (1) used above was of this form. In this special case, the map is monotonic in $\psi$. In this section, we do not make any assumption of this kind. We assume that the $s$ function defined above is bounded below by a positive constant and bounded above by another constant.
It can be seen that the induction of Section 3 extends without modifications to this more general setting. Similarly, blocks can be defined in the same way as in Subsection 4.3. Using this, one can show that the stability condition of this general class of systems is also of the form
${\mathbb{E}}T>{\mathbb{E}}\widehat{\sigma}$
(with the notation of Section 4), with the same interpretation of $\widehat{\sigma}$ as the time to serve a block when all elements of the said block are available at time zero and are served in the FCFS order prescribed by the constraints at the speed prescribed by the interaction function $s$.
7 Discussion and Open Questions
7.1 On the complexity of this class of systems
The first set of questions concerns the complexity of the estimation of $\lambda_{c}$ and of other statistical properties of this class of dynamical systems. If there are a few cases where one can get closed-form expressions for $\lambda_{c}$, even the simulation of a perfect sample of the two random variables allowing one to evaluate $\lambda_{c}$ (see Subsection 5.1) may be a challenging problem. This is well illustrated by the following discrete model: the loci of arrivals are uniformly distributed on the integers ranging from 1 to $N$ and the exclusion radii are equal to $2$ (namely a customer of locus $k$ can only be served after all customers of locus $k-2,k-1,k,k+1$ and $k+2$ that arrived before have left - conditions involving integers not in $[1,N]$ are void). Assuming that $N$ is even, a natural collection of sets implementing the conditions of Subsection 4.1 is $B_{1}$ the ball (of diameter 3) centered in 2, $B_{2}$ that centered in 4, etc, up to $B_{K}$ with $K=\frac{N-1}{2}$ that centered in $N-1$. Then the probability of event $C$ is
$\left(\frac{1}{N}\right)^{K}.$
For $N=21$, the mean number of arrivals until event $C$ is $21^{10}>10^{13}$.
7.2 Connection to the Saturation Rule
The results obtained in Section 6 can be seen as a partial extension of the saturation rule of [11] to nonmonotonic dynamics (in contrast to PHHG for instance). For all models of spatial interaction queues considered there, the stability region is obtained when assessing the time to empty the system when blocking arrivals, exactly as in the saturation rule.
7.3 Open problems
We now list a few open questions.
  1. 1.
    The saturation rule was established in the stationary ergodic framework. In contrast, the results of the present paper require independent assumptions. Can one relax some of these assumptions and still be able to build the stationary regime?
  2. 2.
    The Poisson Hail model was successfully defined on the infinite Euclidean space. Can one construct the PHWG model and extend the stability results obtained in the present paper to non-compact phase spaces? On most sensible models, there will be no zigzag event of positive probability. Therefore, another mathematical approach will be needed.
  3. 3.
    Can one get bounds or estimates on the stationary system time of a customer arriving at a given locus, on the stationary point process of customers in service, or on the total number of customers waiting?
  4. 4.
    The basic dynamics is fundamentally nonmonotonic in that delaying the arrival of a given customer may lead to an earlier departure for another customer. This is why the saturation rule cannot be applied. Can one circumvent this difficulty, e.g., by averaging over permutations of the exclusion sets?
  5. 5.
    It would be practically useful to adapt the model described here to a situation closer to the CSMA/CA setting. This will in particular require the extension of this model to disciplines different from FCFS and more precisely to disciplines based on the carrier sensing and back-off mechanisms used in this context. Conversely, it would be interesting to devise concrete distributed algorithms that allow one to implement the model described in the present paper.
Acknowledgments
The work of F. Baccelli was supported by the European Research Council project titled NEMO under grant ERC 788851 and by the French National Agency for Research project titled France 2030 PEPR réseaux du Futur under grant ANR-22-PEFT-0010.
8 Appendix
8.1 Discrete time
In this subsection, we reformulate our Theorem 1 in a simpler and more general form, as Proposition 13, and provide its proof.
Assume that, for each $x\geq 0$, there is given an i.i.d. sequences $\{\xi_{n}(x)\}$ of random variables with finite mean ${\mathbb{E}}\xi_{1}(x)$, and that all these sequences are mutually independent. Introduce the following recurrence relation: $X_{0}=const$ a.s. and
$X_{n+1}=(X_{n}+\xi_{n}(X_{n}))^{+},\ \ n\geq 1.$ (16)
The sequence $\{X_{n}\}$ is a time-homogeneous Markov chain, and the following result holds.
Proposition 13.
Assume that the family $\{\xi_{1}(x)\}_{x\geq 0}$ is uniformly integrable and that
$\xi_{1}(x)\rightarrow_{st}\xi\equiv\xi(\infty),\ \text{ as }\ x\to\infty.$ (17)
  1. (i)
    If ${\mathbb{E}}\xi<0$, then there exists a finite $K>0$ such that the set $[0,K]$ is positive recurrent, i.e., $\tau_{x}\equiv\tau_{x}(K)=\min\{n\geq 1:X_{n}\leq K\}<\infty$ a.s., for any $X_{0}=x\geq 0$ and, moreover,
    $\sup_{0\leq x\leq K}{\mathbb{E}}\tau_{x}<\infty.$ (18)
    If, in addition, the Markov chain has a positive recurrent atom at point, say, $x_{0}$ and if ${\mathbb{P}}(X_{1}=x_{0}\ |\ X_{0}=x_{0})>0$, then there exist a unique stationary distribution, say $\pi$, and
    ${\mathbb{P}}(X_{n}\in\cdot\ |X_{0}=x)\to\pi(\cdot),\text{as}\ n\to\infty,$ (19)
    in the total variation, for any initial value $X_{0}=x$.
  2. (ii)
    If ${\mathbb{E}}\xi_{1}>0$ and if there is a constant $\gamma>0$ such that, for any $x$,
    ${\mathbb{P}}(\xi_{1}(x)>\gamma)\geq\gamma,$ (20)
    then, for any initial $x\geq 0$, we have
[…]
    and, in particular, $X_{n}\to\infty$ a.s. as $n\to\infty$.
Remark 14.
In the case where ${\mathbb{E}}\xi=0$, there may be various scenarios, including positive recurrence, null recurrence and transience – depending on the way the distribution of $\xi_{1}(x)$ converges to the distribution of $\xi$ as $x$ grows.
Before providing the proof of Proposition 13, we recall the classical Foster criterion and two of its corollaries.
Proposition 15.
Consider a time-homogeneous Markov chain $\{X_{n}^{(x)}\}$ on a measurable state space $({\cal X,B_{X}})$ that starts from initial state $X_{0}^{(x)}=x$. Let $L:{\cal X}\to\infty$ be an unbounded measurable function. Let
$\delta_{x}=L(X_{1}^{(x)})-L(x).$
Assume there exists $\varepsilon>0$ and $K>0$ such that
$\sup_{y:L(y)\leq K}{\mathbb{E}}\delta_{y}<\infty\quad\text{and}\quad{\mathbb{E}}\delta_{x}\leq-\varepsilon,\ \ \text{if}\ \ L(x)>K.$
Then the set $B:=\{x:\ L(x)\leq K\}$ is positive recurrent.
The following two corollaries are straightforward. The proof of the first one may be found, e.g., in Chapter 6 of [12].
Corollary 16.
Let $D$ be any subset of the set $B$ of the last proposition. Denote $\psi_{x}=\min\{n\geq 1:\ X_{n}^{(x)}\in D\}$. If, in the conditions of the Foster criterion, we have in addition that there exist $n_{0}\geq 1$ and $p>0$ such that
${\mathbb{P}}(\psi_{x}\leq n_{0})\geq p,\ \ \text{for all}\ \ x\in B,$
then the set $D$ is positive recurrent, too.
Corollary 17.
Under the conditions of Corollary 16, assume in addition that the set $D$ consists of a single point and that the Markov chain is aperiodic. Then the Markov chain admits a single stationary distribution, say $\pi$, and, for any initial state $x\in{\cal X}$, the distribution of the Markov chain $X_{n}^{(x)}$ converges to $\pi$ in the total variation norm.
We are now in a position to provide the proof of part (i) of Proposition 13.
Proof.
We let $\xi_{n}(x)=\sigma_{n}(x)-t_{n}$. Further, let $\varepsilon:=-{\mathbb{E}}\xi>0$ and $\Delta(x)={\mathbb{E}}(X_{1}-X_{0}\ |\ X_{0}=x)$. We have
$\sup_{x\geq 0}\Delta(x)<\infty\quad\text{and}\quad\lim_{x\to\infty}\Delta(x)=-\varepsilon<0.$
Therefore, there exists $K>0$ such that
$\sup_{x\in[0,K]}\Delta(x)<\infty\quad\text{and}\quad\sup_{x>K}\Delta(x)\leq-\varepsilon/2<0.$
Then the Foster criterion can be applied.
Further, the existence of a positive recurrent atom and aperiodicity imply the stability result.
In the particular case of Theorem 1, the two latter properties are guaranteed by unboundedness of the support of the distribution of interarrival times.
Before giving the proof of the statement (ii) of Proposition 13, we recall Theorem 35 in Section 11 of [12] which is formulated below. Its proof is a minor modification of the proof of the corresponding result from PhD thesis by [13]. and, in turn, the latter result is a slight generalization of the instability theorem in [14].
Consider a time-homogeneous Markov chain $\{X_{n}^{(x)}\}$ on a measurable state space $({\cal X,B_{X}})$ that starts from initial state $X_{0}^{(x)}=x$. Let $L:{\cal X}\to\infty$ be an unbounded measurable function. For $K>0$, let
$\tau_{x}(K)=\min\{n\geq 1\ :\ X_{n}^{(x)}>K\}.$
Let
$\delta_{x}=L(X_{1}^{(x)})-L(x).$
The result of Theorem 35 from [12] may be formulated as follows
Theorem 3.
Assume there exist numbers $\varepsilon>0,K>0,M>0$ and an integrable random variable $\eta>0$ a.s. such that
  1. 1.
    $\tau_{x}(K)<\infty$ a.s., for all $x\in{\cal X}$;
  2. 2.
    for all $x\in{\cal X}$ such that $L(x)>K$,
    ${\mathbb{E}}\min(\delta_{x},M)\geq\varepsilon$ (21)
    and
    $\min(\delta_{x},0)\geq_{st}-\eta.$ (22)
Then, for any $x\in{\cal X}$,
[…]
and, therefore,
${\mathbb{P}}\left(\lim_{n\to\infty}L(X_{n}^{(x)})=\infty\right)=1.$
We are now in a position to prove the second statement of Proposition 13.
Proof.
Here $({\cal X,B_{X}})$ is the positive half-line with the Borel sigma-algebra and $L(x)=x$. We again let $\xi_{n}(x)=\sigma_{n}(x)-t_{n}$. Then condition 1 of the theorem follows from condition (13), condition (21) from the uniform integrability of $\{\xi_{1}(x),x\geq 0\}$, and condition (17), and condition (22) follow from the fact that the random variables $\{t_{n}\}$ are i.i.d. with a finite mean, thus one can take $\eta=t_{1}$. ∎
8.2 Continuous time, proof of Theorem 2
Proof of (i). This follows from known results for regenerative processes – see e.g. Theorem 6.1.2 and Proposition 7.3.8 in [15].
Proof of (ii). This follows directly from the following observations: First,
$W(t)\geq W_{\eta(t)}-Z_{\eta(t)}\quad\text{for any}\quad t\geq 0,$
where $\eta(t)=\min\{n:T_{n}\geq t\}$ and $\{Z_{m}\}$ are defined in (10). Since $Z_{m}$ are i.i.d. with a finite mean and since $\eta(t)\to\infty$ a.s.,
$\frac{W(t)}{t}\equiv\frac{W(t)}{\eta(t)}\cdot\frac{\eta(t)}{t}\geq\frac{W_{\eta(t)}}{\eta(t)}\cdot\frac{\eta(t)}{t}-\frac{\eta(t)}{t}\cdot\frac{\max_{1\leq i\leq\eta(t)}Z_{i}}{\eta(t)}$
and, by the Strong Law of Large Numbers,
$\liminf_{t\to\infty}\frac{W(t)}{t}\geq\varepsilon\cdot\frac{1}{{\mathbb{E}}(T_{2}-T_{1})}>0,$
since
$\frac{\max_{1\leq i\leq n}Z_{m}}{m}\to 0\quad\text{a.s., as}\quad m\to\infty.$